OpenStudy (vortish):

make a scatter plot using the given points (5,7) (-1,6)(0,-6) (-3,2)

OpenStudy (zepp):

You want to find the slope that represents the best the situation right?

OpenStudy (vortish):

yes it give multiple choices for the situtation zeep

OpenStudy (zepp):

Ok good, are you familiar with Mayer's method?

OpenStudy (vortish):

no

OpenStudy (zepp):

What about the rectangle method (approximation method)?

OpenStudy (vortish):

its greek to me

OpenStudy (zepp):

Okay let me explain Mayer's method, it's really easy easy.

OpenStudy (zepp):

First of all, order your points from the lowest value of x to the highest value of x.

OpenStudy (zepp):

(-3,2) (-1,6) (0,-6) (5,7) Now split it in half so both part contains the same number of points. *splits* (-3,2) (-1,6) _________ (0,-6) (5,7) Voila!

OpenStudy (vortish):

got it zepp

OpenStudy (zepp):

Good!, Now calculate the average of each section \((-3+-1)/2=(-3-1)/2=(-4)/2=-2\) \((6-2)/2=4/2=2\) Average point of the first part is \((-2,2)\) \((5+0)/2=5/2=2.5\) \((7--6)/2=13/2=6.5\) Average point of the second part is \((2.5,6.5)\) Is it understandable?

OpenStudy (zepp):

Now you have two points \((-2,2)\) and \((2.5,6.5)\) and you are asked to find the slope, you know what to next :)